OpenStudy (anonymous):

if current drops 1% decrease in power will be?

OpenStudy (anonymous):

among these which one is correct 1%,2%,4%,0.5%?

OpenStudy (chenping):

haha, where's your constant? without a constant, it's impossible to calculate it. Let's assume that Voltage is constant, then P=VI and P1/P2=I1/I2 (proportional). So as the qs stated that I2 is 101% of I1, so power would also recede by 1 % if current is droped by 1%

OpenStudy (anonymous):

P is the power when current is I and after drop it will be P' so decrease in power will be P-P' So P"=P-P'= 0.01P ie 1%